OpenStudy (anonymous):

Find the values of y and z. 2y+12 4z+20 4y-18

OpenStudy (anonymous):

Please help I'll fan and medal.

OpenStudy (anonymous):

Wut

OpenStudy (anonymous):

I dont understand what I'm supposed to to.

OpenStudy (anonymous):

So basically it's 2y+12=4y-18=4z+20, yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well then separate it like this ( 2y+12=4y-18) = 4z+20. Now solve for y first.

OpenStudy (anonymous):

oh ok now i think i get it, thanks.

OpenStudy (anonymous):

ye no problem (: